The state of Wyoming is rectangular shape. The length of this rectangular state is 185 miles less than twice the width. If the state's perimeter is 1280 miles, what are the dimensions?
Answer by checkley71(8403)

L=2W-185
2(2W-185)+2W=1280
4W-370+2W=1280
6W=1280+370
6W=1650
W=1650/6
W=275 MILES IS THE WIDTH.
L=2*275-185
L=550-185
L=365 IS THE LENGTH.
PROOF
2*275+2*365=1280
550+730=1280
1280=1280